OpenStudy (anonymous):

Which of the following numbers can be expressed as repeating decimals? 2/9, 3/8, 5/6, 5/4 3/8 and 5/6 2/9 and 5/6 2/9 and 5/4 3/8 and 5/4

OpenStudy (anonymous):

@Mr.Z

OpenStudy (mr.z):

Excuse me?... Those all look like whole numbers to me. What math class is this for?

OpenStudy (anonymous):

@mr.z these are fractions. 3 over 8 would mean 3/8

OpenStudy (mr.z):

B in my opinion.

OpenStudy (mr.z):

Not B!

OpenStudy (mr.z):

3/8 and 5/6 This one I think.

OpenStudy (anonymous):

@mr.z any reason? this is really important o3o

OpenStudy (mr.z):

3/8 = 0.375 5/6 = 0.83333333333 Keep on repeating those and I don't believe you'll get a whole number. Or maybe it's 0.22222222222 (2/9) Yes, it's B. 2/9 = 0.22222222222 5/6 = 0.83333333333